3 Direct Materials and Direct Labor Variances At the beginning of June, Kimber Toy Company budgeted...

3

Direct Materials and Direct Labor Variances

At the beginning of June, Kimber Toy Company budgeted 14,000 toy action figures to be manufactured in June at standard direct materials and direct labor costs as follows:

Direct materials $14,000
Direct labor 13,720
Total $27,720

The standard materials price is $0.50 per pound. The standard direct labor rate is $14.00 per hour. At the end of June, the actual direct materials and direct labor costs were as follows:

Actual direct materials $13,300
Actual direct labor 13,000
Total $26,300

There were no direct materials price or direct labor rate variances for June. In addition, assume no changes in the direct materials inventory balances in June. Kimber Toy Company actually produced 12,900 units during June.

Determine the direct materials quantity and direct labor time variances. Round your per unit computations to two decimal places, if required. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct materials quantity variance $
Direct labor time variance $
Answer #1

When there is no direct material price and direct labor rate variance then total variance is equal to material quantity and time variance

Direct material quantity variance = (14000/14000*12900)-13300 = 400 Unfavorable

Direct labor time variance = (13720/14000*12900)-13000 = 358 Unfavorable